OpenStudy (sierra!):

Kayla designed a rectangular patio that was 3 times as long as it was wide and had a perimeter of 160 feet. What was the area of the patio? A. 600 ft2 B. 1200 ft2 C. 1800 ft2 D. 2400 ft2

OpenStudy (sierra!):

@perl

OpenStudy (sierra!):

@Data_LG2

OpenStudy (sierra!):

@iGreen

OpenStudy (sierra!):

@XoALIEoX

undeadknight26 (undeadknight26):

P = 2L + 2W L = 3W (because it's 3 times longer than it is wide) Sub L into your expression for P to solve for W, P = 2L + 2W P = 2(3W) + 2W P = 8W W = P / 8 W = 160 ft / 8 W = 20 ft Since you have W, you can solve for L, L = 3W L = 3(20 ft) L = 60 ft Now you can find the area, A = LW A = (60 ft)(20 ft) A = 1200 ft² Taken from: https://answers.yahoo.com/question/index?qid=20130606183730AAXgDv8

OpenStudy (sierra!):

Thank you!

OpenStudy (anonymous):

^^^ @undeadknight26 explained it and showed you how to XD beat me to it

undeadknight26 (undeadknight26):

Any time dawg #GoogleFtw